What is the distance between points (5,-3) and (0,9) in a coordinate system?

Answer by checkley77(12569)

(5,-3) and (0,9)
You have a right triangle with sides of x & y & an hypotenuse (distance) of c.
x=5-0=5
y=9+3=12
x^2+y^2=c^2
5^2+12^2=c^2
25+144=c^2
169=c^2
c=sqrt169
c=13 ans. for the distance between these 2 points.